OpenStudy (anonymous):

frank has devised a formula for his catering business that calculates the number of meatballs he needs to prepare. how many meat balls are required for a party of 10 adults and 5 children? please help !

OpenStudy (btaylor):

Do you have the formula?

OpenStudy (anonymous):

the formula is 4a+2c c= number of children m= number of meatballs a= number of adults

OpenStudy (btaylor):

It looks like you are given a (10) and c (5), so just plug those values into the formula m=4a+2c n=4(10)+2(5) m=40+10 m=50 meatballs

OpenStudy (anonymous):

thank you sooo much !! :)

OpenStudy (btaylor):

you're welcome!